# Infra-red thermographic inversion in ST40

### v1

@@ Introduction
Following from the dimensionality of which in Sec. 3.1.2, and the unsteady Eq. (2), FAHF is configured as a 2D time-dependent heat conduction solver in cartesian coordinates ($xy$ plane). The finite difference (FD) method is relied upon for the spatial discretisation in the nodes of the rectangular domain, for reasons of simplicity and speed.
In particular, $N_{x}$ points $\{x_{i}\}_{i=1}^{N_{{x}}}$ are placed along each sub-chord ($x$ direction), with uniform spacing $\Delta x=x_{i+1}-x_{i}\;\forall i\in[1;N_{x})$. Instead, $N_{y}$ points $\{y_{j}\}_{j=1}^{N_{{y}}}$ are distributed along the tile depth ($y$ direction), with uniform spacing $\Delta y=y_{j+1}-y_{j}\;\forall j\in[1;N_{y})$ (potentially differing from $\Delta x$), where $j=1$ identifies the plasma-facing surface, and $j=N_{y}$ the back-surface. The second-order centred FD scheme[12] is then adopted in the interior nodes $(i,j)\in[2;N_{x}-1]\times[2;N_{y}-1]$.
In time, the explicit first-order forward Euler time-marching scheme is implemented at the time instants $\{t^{n}\}_{n=1}^{N_{t}}\in[0;200]$ ms, with uniform time-step $\Delta t=t^{n}-t^{n-1}\;\forall n\in(1;N_{t}]$. This implementation must satisfy the stability criterion [12]:
$\Delta t\cdot\left(\frac{1}{\Delta x^{2}}+\frac{1}{\Delta y^{2}}\right)<\frac{1}{2\alpha}.$ (3)
@@
At the plasma-facing boundary of the computational domain ($j=1$), a Dirichlet-type boundary condition (BC) is enforced via the experimental temperature data. Exploiting the Calcam software[47] according to Appendix A, $T_{\text{IR}}(x,\phi)$ is sampled over the set $\{x_{i}\}_{i=1}^{N_{x}}$ along each chord at every time instant $\{t^{n}\}_{n=1}^{N_{\text{t}}}$, obtaining the BC data $T_{i,1}^{n}=T_{\text{IR},i}^{n}\;\forall i,n$. Although both the user-selected $\Delta x$ and the $\Delta t$ resulting form the stability criterion (Eq. (3)) might fall below the IR camera spatial and/or temporal resolution (Sec. 2.2), this is not here of concern, as the linear interpolation of $T_{\text{IR}}$ is checked to not introduce any artifacts (e.g. over-/under-shooting).
On the remaining non-plasma-facing boundaries ($\{i=1\}\cup\{i=N_{x}\}\;\forall j>1$ or $j=N_{y}\;\forall i$), first-order homogeneous Neumann BCs are enforced. Physically, homogeneous Neumann BCs represent adiabatic (no-flux) boundaries, and implicitly assume the sub-chords (divertor tiles) to be thermally independent ($\{i=1\}\cup\{i=N_{x}\}\;\forall j>1$). This holds under the legitimate assumptions of a negligible black-body emission in the gap ($T_{\text{IR}}<50$), and of a negligible heat convection/conduction provided by the rarefied neutral gas filling this volume (few Pa in pressure at most [50]). Additionally, the thermal contact of the tile back-surface ($j=N_{y}\;\forall i$) with the supporting structures is also neglected, according to Sec. 3.1.1 -- the variation of the bulk temperature anywhere more than few millimetres away from the plasma-facing boundary being negligible.
In the temporal domain, the initial condition (IC, $t=0$ ms) $T_{i,j}^{1}$ chosen at the first time instant $n=1$ is a uniform distribution in the entire domain. Its value is the spatial-average temperature along the sub-chord as measured by the IR camera, i.e. $T_{i,j}^{1}=\sum_{i=1}^{N_{x}}T_{\text{IR},i}^{1}/N_{x}\;\forall i,j$. Different flavours of such IC are tested, not providing any appreciable difference in the outcome after a few time-steps.
@@
$\begin{split}&\begin{split}\frac{T_{i,j}^{n}-T_{i,j}^{n-1}}{\Delta t}&=\alpha\cdot\frac{T_{i+1,j}^{n-1}-2T_{i,j}^{n-1}+T_{i-1,j}^{n-1}}{\Delta x^{2}}+\\ &+\alpha\cdot\frac{T_{i,j+1}^{n-1}-2T_{i,j}^{n-1}+T_{i,j-1}^{n-1}}{\Delta y^{2}}+\\ &+o(\Delta x^{2})+o(\Delta y^{2})+o(\Delta t).\end{split}\\ \\ &\begin{split}&\text{IC: }T^{1}_{i,j}=\sum_{i=1}^{N_{x}}T_{\text{IR},i}^{1}/N_{x}\text{ if $n=1\;\forall i,j$.}\\ &\text{BC: }\text{Dirichlet }T^{n}_{i,1}=T_{\text{IR},i}^{n}\text{ if $j=1\;\forall i,n\geq 2$.}\\ &\begin{split}\text{BC: }&\text{homogeneous Neumann (no-flux)}\\ &\text{if $\{i=1\}\cup\{i=N_{x}\}\;\forall j>1,n\geq 2$}\\ &\text{or $j=N_{y}\;\forall i,n\geq 2$.}\end{split}\end{split}\end{split}$ (4)
The only unknown $T_{i,j}^{n}\;\forall n\in[2;N_{t}]$ is explicitly retrieved at each timestep, and embodies the output of the simulation with a discretisation error $o(\Delta x^{2})+o(\Delta y^{2})+o(\Delta t)$.
Surface heat flux and Eich fit
@@
The quantity of primary physical interest of the IR thermographic inversion is the surface heat flux $q\;(\text{W}\cdot\text{m}^{-2})$ perpendicular to the PFC surface (i.e. along $y$). From the temperature distribution $T_{i,j}^{n}$ computed by the algorithm via Eq. (4), the surface heat flux is computed in post-processing at the plasma-facing surface ($j=1$) according to the first-order discretisation:
$\begin{split}q_{i}^{n}=&-\kappa\cdot\frac{T_{i,2}^{n}-T_{i,1}^{n}}{\Delta y}+\\ &+o(\Delta x^{2})+o(\Delta y)+o(\Delta t),\end{split}$ (5)
@@
with $q_{i}^{n}=q_{i,1}^{n}$. The error $o(\Delta x^{2})+o(\Delta t)$ is inherited from $T_{i,j}^{n}$ of Eq. (4), while $o(\Delta y^{2})\ll o(\Delta y)$ by definition. Higher-order spatial discretisations[49] are available, but here not required.
Once $q_{i}^{n}$ is known for two adjacent sub-chords, fitting of the juxtaposed profile (Sec. 3.1.2) is accomplished by leveraging the Eich function $q_{\text{Eich}}$ from Eq. (1) of Ref. [6]. Although an iterative procedure is actually best-posed in presence of the multi-parameter non-linear fitting function[46] $q_{\text{Eich}}$, this is not here accomplished for the sake of simplicity. Instead, the fit is deemed satisfactory upon showing $R^{2}>98.5\%$. This simplified approach still allows to convey an equivalent message.
Ultimately, the main output of the fitting procedure is embodied by the physical parameters of: maximum heat flux $q_{0}$, power fall-off length $\lambda_{q}$, spreading factor $S$, background heat flux $q_{\text{BG}}$, strike point location $x_{0}$, and poloidal flux expansion $f_{X}$. The reader is redirected to Ref. [6] for detailed descriptions of such quantities. The integral value of the Eich profile $q_{\text{Eich}}(x)$ is here additionally added, and defined as:
[…]
Internal energy variation and incoming energy
Sections 3.1.5 and 3.1.6 provide the building blocks to compute the internal energy variation (per unit length, along the non-modelled $\phi$ direction) throughout the volume of the domain:
$\begin{split}\Delta\mathcal{E}_{\text{int}}^{n}&=\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{y}}\varrho c_{\text{p}}(T_{i,j}^{n}-T_{i,j}^{n-1})\cdot\Delta x\Delta y+\\ &+o(\Delta x^{2})+o(\Delta y^{2})+o(\Delta t),\end{split}$ (7)
and the net energy (per unit length) flowing through the plasma facing boundary:
$\begin{split}\mathcal{E}_{\text{surf}}^{n}&=\sum_{i=1}^{N_{x}}q_{i}^{n}\cdot\Delta x\Delta t+\\ &+o(\Delta x^{2})+o(\Delta y)+o(\Delta t),\end{split}$ (8)
which enable the characterisation of the energy-preserving properties of FAHF. In Eqs. (7) and (8) the errors follow from Eqs. (4) and (5), respectively.
@@
COMSOL Multiphysics® [42], herein abbreviated to ”COMSOL”, is a widely-adopted state-of-the-art tool for multiphysical modelling. This suite is here configured as a pure heat conduction solver, and therefore utilises the finite element method (FEM) in space, with quadratic Lagrange elements, and second-order backward time stepping [42].
A variety of different models, detailed in Sections 4.2.1 and 4.2.2, is considered in COMSOL. This is an attempt of matching FAHF the closest in dedicated instances, and quantifying the model/geometry errors of FAHF in others, ultimately resulting in an estimate of its accuracy.
The same computational domain of FAHF (Sec. 3.1.2) is uniformly meshed with $N_{x}\times N_{y}=100\times 100$ rectangular elements ($\Delta x=3.5\cdot 10^{-4}$ m and $\Delta y=2.9\cdot 10^{-4}$ m), and the same initial and boundary conditions (Sec. 3.1.4) are adopted, with $\Delta t=1\cdot 10^{-4}$ s. Because of reasons related to ingent computational cost, only one divertor tile (bottom left in Fig. 1b) is modelled by COMSOL in 3D and, for consistency, only the corresponding sub-chord of chord 10 (Sec. 3.3.1) belonging to such tile is modelled in 2D. Clearly, this time-saving choice does not impact any of the conclusions inferred.
Ancillary convergence tests of the COMSOL results have been performed, but are not here reported. Noteworthy is instead the need of de-activating the automatic smoothing of the resulting heat flux in COMSOL, which would otherwise constitute a further unwanted source of error, for not being implemented in FAHF.
@@
Reference chord
Chord number 10 ($k=10$, Sec. 3.1.2) is taken as a reference, and depicted in Fig. 1b (blue dots). This chord sits sufficiently far away from poloidal-running gaps to not show any deleterious boundary effect [46]. Also, chord 10 is affected neither by magnetic shadowing [46] nor by surface imperfections (Sec. 3.1.2), both which would spuriously distort the heat flux pattern.
Throughout the following, only chord 10 is investigated, but similar results apply for all the other ones. The output of a representative multi-chord analysis is reported by Ref. [46].
@@
The behaviour of the errors in time $\varepsilon_{*}(\Delta t)$ and space $\varepsilon_{*}(\Delta x)$ is assessed via time and space convergence, the latter forcing $\Delta x=\Delta y$ for the sake of simplicity. These procedures are a numerical characterisation of the precision of the algorithm (verification), ensure trust in the internal workings of the code, and suggest the most suitable choice of $\Delta t$ and $\Delta x$, within the limit posed by Eq. (3). These processes only involve the actual output of the simulation, i.e. the temperature distribution $T_{i,j}^{n}$ (Sec. 3.1.5), and are agnostic of the heat flux $q_{i}^{n}$ (Sec. 3.1.6).
In particular, in the time convergence, the time-step $\Delta t$ is varied in the interval $[4\cdot 10^{-6};4\cdot 10^{-4}]$ s. The actual values are detailed in Table 1. Instead, Table 2 reports the values $\Delta x\in[6.8\cdot 10^{-5};2.3\cdot 10^{-3}]$ m of the space convergence. Linear interpolation is employed whenever $\Delta t<\Delta t_{\text{IR}}=1.25\cdot 10^{-3}$ s or $\Delta x<\Delta x_{\text{IR}}=2.1\cdot 10^{-4}$ m (Sec. 2.2).
Each simulation in the time convergence is run with a fixed spatial discretisation with $\Delta x=5.5\cdot 10^{-4}$ m. This space-step is small enough such that to not alter the error in time across all the cases, i.e. $o(\Delta x^{2})\ll o(\Delta t)\;\forall\Delta t$ in the time convergence. In the space convergence, $\Delta t=7.2\cdot 10^{-6}$ s dually guarantees $o(\Delta t)\ll o(\Delta x^{2})\;\forall\Delta x$.
Common to both procedures, a reference point in time $t_{*}=128$ ms and a reference point in space $(x_{*},y_{*})=(16.55;2.21)$ mm are selected where to locally compare the temperature across the various discretisations.
The time point $t_{*}=128$ ms lies well within $[90;140]\;\text{ms}$, the time window of most interest (Sec. 2.3). Unless otherwise specified, the results obtained at $t=t_{*}$ apply to any other time instant. The point $(x_{*},y_{*})$ lies in the volume of the domain, within $\sim 3$ mm of the plasma-facing surface so as to be appreciably affected by a temperature variation (Sec. 3.1.1), and away from any other boundaries, making it suitable for comparison across different cases. This ”local” approach is preferred over integral errors (in the fashion of Sec. 3.3.3) to minimise any near-boundary artifacts which would impair a precise convergence analysis.
The reference temperature $\tilde{T}_{*}=\tilde{T}(x_{*},y_{*},t_{*})$ to compare the other cases against is defined as the result obtained with the smallest time/space-step of the time/space convergence, i.e. the most precise instance. The relative difference of the resulting $T_{*}=T(x_{*},y_{*},t_{*})$ of a given simulation with respect to the reference is labelled as ”error” and it is computed according to:
$\varepsilon_{*}=\frac{\left|T_{*}-\tilde{T}_{*}\right|}{\left|\tilde{T}_{*}\right|}.$ (9)
Clearly, the error of the very reference case is zero by definition, and it is therefore discarded.
@@
With regards to the surface heat flux (Sec. 3.1.6), the relative difference of the result $q_{i}^{n}$ of a given simulation (e.g. COMSOL, Sec. 3.2) with respect to FAHF’s $\tilde{q}_{i}^{n}$ is referred to as ”heat flux error” $\varepsilon_{q}$. Such error leverages the $\ell^{1}$ norm, with the heat fluxes being interpolated in time and space onto the grid of the most refined case between the two.
@@
$\varepsilon^{n}(q)=\frac{\sum_{i=1}^{N_{x}}\left|q_{i}^{n}-\tilde{q}_{i}^{n}\right|}{\sum_{x=1}^{N_{x}}\left|\tilde{q}_{i}^{n}\right|},$ (10)
@@
$\varepsilon_{i}(q)=\frac{\sum_{n=1}^{N_{t}}\left|q_{i}^{n}-\tilde{q}_{i}^{n}\right|}{\sum_{n=1}^{N_{t}}\left|\tilde{q}_{i}^{n}\right|},$ (11)
returning an error profile in space along $x$ (at the plasma-facing surface). This ”integral” approach allows accounting for all the areas of agreement/disagreement among two cases.
@@
For a perfect energy balance to be satisfied, the internal energy variation (Eq. (7)) must match the net incoming energy (Eq. (8)). The relative difference among $\Delta\mathcal{E}_{\text{int*}}=\Delta\mathcal{E}_{\text{int}}(t_{*})$ and $\mathcal{E}_{\text{surf*}}=\mathcal{E}_{\text{surf}}(t_{*})$ for each case, labelled as ”energy balance error”, is thus defined by:
$\varepsilon_{*}(\mathcal{E})=\frac{\left|\Delta\mathcal{E}_{\text{int*}}-\mathcal{E}_{\text{surf*}}\right|}{\left|\Delta\mathcal{E}_{\text{int*}}+\mathcal{E}_{\text{surf*}}\right|/2}.$ (12)

### v2

@@ Introduction
The finite difference (FD) method is relied upon for the 2D spatial discretisation in the nodes of the rectangular domain, for reasons of simplicity and speed.
In particular, $N_{x}$ points $\{x_{i}\}_{i=1}^{N_{{x}}}$ are placed along each sub-chord ($x$ direction), with uniform spacing $\Delta x=x_{i+1}-x_{i}\;\forall i\in[1;N_{x})$. Instead, $N_{y}$ points $\{y_{j}\}_{j=1}^{N_{{y}}}$ are distributed along the tile depth ($y$ direction), with uniform spacing $\Delta y=y_{j+1}-y_{j}\;\forall j\in[1;N_{y})$ (potentially differing from $\Delta x$), where $j=1$ identifies the plasma-facing surface, and $j=N_{y}$ the back-surface. The second-order centred FD scheme[10] is then leveraged in the interior nodes $(i,j)\in[2;N_{x}-1]\times[2;N_{y}-1]$.
Non-uniform grids along $y$ are found in the literature [18], but are advantageous under specific circumstances outside the scope of the present work (e.g. when including an impurity surface layer[59]). The space convergence analysis of Sec. 5.1.1 justifies the adoption of a uniform mesh.
In time, the first-order explicit forward Euler time stepping scheme is implemented at the time instants $\{t^{n}\}_{n=1}^{N_{t}}\in[0;200]$ ms, with uniform time-step $\Delta t=t^{n}-t^{n-1}\;\forall n\in(1;N_{t}]$. This implementation must satisfy the stability criterion [10]:
$\Delta t\times\left(\frac{1}{\Delta x^{2}}+\frac{1}{\Delta y^{2}}\right)<\frac{1}{2\alpha}\>.$ (7)
Reasons behind selecting an explicit scheme are the ease of implementation and, with the proper user settings identified in Sec. 5.1.1, a satisfactory speed. An implicit implementation would be considered, were it needed for specific applications[18] beyond those of the present work.
@@
With the BC choices of Sec. 3.1, and exploiting the Calcam software[52] according to Appendix Calcam, the experimental datum $T_{\text{IR}}(x,\phi)$ is sampled over the set $\{x_{i}\}_{i=1}^{N_{x}}$ along each chord at every time instant $\{t^{n}\}_{n=1}^{N_{\text{t}}}$. The Dirichelet BC data $T_{i,1}^{n}=T_{\text{IR},i}^{n}\;\forall i,n$ is obtained, and enforced at the plasma-facing boundary ($j=1$). Although both the user-selected $\Delta x$ and the $\Delta t$ resulting form the stability criterion (Eq. (7)) might fall below the IR camera spatial and/or temporal resolution (Sec. 2.2.1), this is not here of concern, as the linear interpolation of $T_{\text{IR}}$ is checked to not introduce any artifacts (e.g. over-/under-shooting). An assessment of the uncertainty on $T_{\text{IR}}$ is provided in Appendix Experimental uncertainty.
A first-order homogeneous Neumann BC is adopted for the non-plasma-facing boundaries ($\{i=1\}\cup\{i=N_{x}\}\;\forall j>1$ or $j=N_{y}\;\forall i$), assumed adiabatic (Sec. 3.1).
In the temporal domain, the IC $T_{i,j}^{1}$ chosen at the first time instant $n=1$ is a uniform distribution in the entire domain. Its value is the spatial-average temperature along the sub-chord as measured by the IR camera, i.e. $T_{i,j}^{1}=\sum_{i=1}^{N_{x}}T_{\text{IR},i}^{1}/N_{x}\;\forall i,j$. Different flavours of such IC are tested, not providing any appreciable difference in the outcome after a few time-steps.
@@
$\begin{split}&\begin{split}\frac{T_{i,j}^{n}-T_{i,j}^{n-1}}{\Delta t}&=\alpha\>\frac{T_{i+1,j}^{n-1}-2T_{i,j}^{n-1}+T_{i-1,j}^{n-1}}{\Delta x^{2}}+\\ &+\alpha\>\frac{T_{i,j+1}^{n-1}-2T_{i,j}^{n-1}+T_{i,j-1}^{n-1}}{\Delta y^{2}}+\\ &+o(\Delta x^{2})+o(\Delta y^{2})+o(\Delta t)\>;\end{split}\\ \\ &\begin{split}&\text{IC: }T^{1}_{i,j}=\sum_{i=1}^{N_{x}}T_{\text{IR},i}^{1}/N_{x}\text{ if $n=1\;\forall i,j$}\>;\\ &\text{BC: }\text{Dirichlet }T^{n}_{i,1}=T_{\text{IR},i}^{n}\text{ if $j=1\;\forall i,n\geq 2$}\>;\\ &\begin{split}\text{BC: }&\text{homogeneous Neumann (no-flux)}\\ &\text{if $\{i=1\}\cup\{i=N_{x}\}\;\forall j>1,n\geq 2$}\\ &\text{or $j=N_{y}\;\forall i,n\geq 2$}\>.\end{split}\end{split}\end{split}$ (8)
The only unknown $T_{i,j}^{n}\;\forall n\in[2;N_{t}]$ is explicitly retrieved at each timestep, and embodies the output of the simulation. While the mathematical properties of the discretisation error $o(\Delta x^{2})+o(\Delta y^{2})+o(\Delta t)$ are listed in Appendix Algebra of the energy balance error, its behaviour in FAHF is thoroughly scrutinised throughout Sec. 5.1, and shown to negligibly impact the main conclusions of the work.
Surface heat flux discretisation
@@
The surface heat flux of Sec. 3.1 is calculated in post-processing at the plasma-facing surface ($j=1$) according to the first-order discretisation:
$\begin{split}q_{i}^{n}=&-\kappa\>\frac{T_{i,2}^{n}-T_{i,1}^{n}}{\Delta y}+\\ &+o(\Delta x^{2})+o(\Delta y)+o(\Delta t)\>,\end{split}$ (9)
@@
with $q_{i}^{n}=q_{i,1}^{n}$. The error $o(\Delta x^{2})+o(\Delta t)$ is inherited from $T_{i,j}^{n}$ of Eq. (8), while $o(\Delta y^{2})\ll o(\Delta y)$ by definition (Eq. 18). Higher-order spatial discretisations[58, 14] could be implemented, but are here not required, as proven in Secs. 5.1.2 and 5.1.3.
Discretisation of internal energy variation and deposited energy
The discretisation of Eqs. 4 and 5 over one time-step ($\tau=[t;t+\Delta t]$) yields:
$\begin{split}\Delta\mathcal{E}_{\text{int}}^{n}&=\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{y}}\varrho c_{\text{p}}(T_{i,j}^{n}-T_{i,j}^{n-1})\times\Delta x\Delta y+\\ &+o(\Delta x^{2})+o(\Delta y^{2})+o(\Delta t)\>,\end{split}$ (10)
for internal energy variation, and:
$\begin{split}\mathcal{E}_{\text{surf}}^{n}&=\sum_{i=1}^{N_{x}}q_{i}^{n}\times\Delta x\Delta t+\\ &+o(\Delta x^{2})+o(\Delta y)+o(\Delta t)\>,\end{split}$ (11)
for the net deposited energy, respectively. This enables the characterisation of the energy balance in FAHF from a purely numerical standpoint, which ensures further trust in the choices of Secs. 4.1.3 and 4.1.4. In Eqs. (10) and (11) the errors follow from Eqs. (8) and (9), respectively.
@@
Overview and motivation
COMSOL Multiphysics®, herein abbreviated to ”COMSOL”, is a widely-adopted state-of-the-art tool for multiphysical modelling[60]. This suite is here configured as a pure heat conduction solver, and utilises the finite element method (FEM) in space with quadratic Lagrange elements, and second-order backward time stepping [60].
Further detailed in Secs. 5.2.1 and 5.2.3, different geometries (2D and 3D) and physical models (w/ and w/o temperature-dependent properties, w/ and w/o CuCrZr) are employed in COMSOL. This is an attempt of quantifying the model/geometry errors of FAHF, ultimately resulting in an estimate of FAHF’s accuracy within its simplifying Asms. #1, #2 and #3.
Setup and limitations of the comparison
Because of reasons related to ingent computational cost, only one divertor tile (bottom left in Fig. 1b) is modelled by COMSOL in 3D and, for consistency, only the corresponding sub-chord belonging to such tile is modelled in 2D. The rectangular domain matching FAHF’s (Sec. 4.1.2) is uniformly meshed in COMSOL with $N_{x}\times N_{y}=100\times 100$ ($N_{\phi}=100$ in 3D) rectangular elements ($\Delta x=3.5\times 10^{-4}$ m and $\Delta y=2.9\times 10^{-4}$ m), and the same initial and boundary conditions (Sec. 4.1.4) are adopted, with $\Delta t=1\times 10^{-4}$ s. Ancillary convergence tests (not shown) have been performed, and confirm that a negligible discretisation error affects the COMSOL results.
However, the fundamental numerics differing in FAHF and COMSOL (FDM vs. FEM, first- vs. second-order time stepping) constitutes a spurious source of disagreement of unknown magnitude. For this reason, the entire Sec. 6.4 is devoted to an in-depth analysis, to carefully motivate the disagreement of the code predictions.
Noteworthy is also the need of de-activating the automatic smoothing of the resulting heat flux in COMSOL, which would otherwise constitute a further unwanted source of error, for not being implemented in FAHF.
@@
Reference chords
From the chord bundle generated for the multi-2D analysis (Sec. 4.1.2), the two chords CH#1 and CH#2 are taken as references for the assessment of Asm. #3 (Sec. 5.2.1), and are depicted in Fig. 1b (blue dots). These chords sit 1.0 mm and 12.5 mm away from the left tile edge along $\phi$, respectively, and are affected neither by magnetic shadowing [48] nor by surface imperfections (Sec. 4.1.2).
If not specified otherwise, CH#2 is implicitly referred to throughout the paper. The output of a complete multi-chord analysis is reported by Ref.[48].
@@
The behaviour of the errors in time $\varepsilon_{*}(\Delta t)$ and space $\varepsilon_{*}(\Delta x)$ is assessed via time and space convergence, the latter forcing $\Delta x=\Delta y$ for the sake of simplicity. These procedures are a numerical characterisation of the precision of the algorithm (verification), ensure trust in the internal workings of the code, and suggest the most suitable choice of $\Delta t$ and $\Delta x$, within the limit posed by Eq. (7). Specific details of these procedures are reported in Appendix Details of time and space convergence, while Tables 3 and 3 condense the actual $\Delta t$ and $\Delta x$ employed.
Given the reference point $(x_{*},y_{*},t_{*})$ defined in Appendix Details of time and space convergence with $t_{*}=128$ ms, the reference temperature $\tilde{T}_{*}$ at such point is defined as the result obtained with the smallest time/space-step of the time/space convergence, i.e. the most precise instance, and used to compare the other cases against. The relative difference of the resulting $T_{*}=T(x_{*},y_{*},t_{*})$ of a certain simulation with respect to the reference is labelled as ”convergence error” and it is computed according to:
$\varepsilon_{*}=\frac{\left|T_{*}-\tilde{T}_{*}\right|}{\left|\tilde{T}_{*}\right|}\>.$ (12)
@@
The relative difference of the resulting surface heat flux $q_{i}^{n}$ (Sec. 4.1.6) of a given simulation (e.g. COMSOL, Sec. 4.2) with respect to FAHF’s $\tilde{q}_{i}^{n}$ (unless specified otherwise) is referred to as ”heat flux error” $\varepsilon_{q}$. Such error leverages the $\ell^{1}$ norm, with the heat fluxes being interpolated in time and space onto the grid of the most refined case between the two.
@@
$\varepsilon^{n}(q)=\frac{\sum_{i=1}^{N_{x}}\left|q_{i}^{n}-\tilde{q}_{i}^{n}\right|}{\sum_{x=1}^{N_{x}}\left|\tilde{q}_{i}^{n}\right|}\>,$ (13)
@@
$\varepsilon_{i}(q)=\frac{\sum_{n=1}^{N_{t}}\left|q_{i}^{n}-\tilde{q}_{i}^{n}\right|}{\sum_{n=1}^{N_{t}}\left|\tilde{q}_{i}^{n}\right|}\>,$ (14)
returning an error profile in space along $x$ (at the plasma-facing surface).
This ”integral” approach conveniently condenses all the areas of agreement/disagreement among two cases. Its only shortcoming is the over-estimation of the actual error if a one specific time interval or space region is of interest.
@@
For a perfect energy balance to be satisfied, the internal energy variation of Eq. (10) must match the net deposited energy of Eq. (11). The relative difference among $\Delta\mathcal{E}_{\text{int*}}=\Delta\mathcal{E}_{\text{int}}(t_{*})$ and $\mathcal{E}_{\text{surf*}}=\mathcal{E}_{\text{surf}}(t_{*})$ for each case, labelled as ”energy balance error”, is thus defined by:
$\varepsilon_{*}(\mathcal{E})=\frac{\left|\Delta\mathcal{E}_{\text{int*}}-\mathcal{E}_{\text{surf*}}\right|}{\left|\Delta\mathcal{E}_{\text{int*}}+\mathcal{E}_{\text{surf*}}\right|/2}\>.$ (15)